Suppose a survey of 24 people asked how many books they have read in 2018. The results are shown in the table below.

Number of Books Read
4 3 8 15 10 0 5 5
7 6 7 1 8 3 3 1
6 2 0 3 7 12 20 2

Step 3 of 7:

Which is a better measure of center for this data? Why?

Give the 5 Number Summary of this data.

What is the interquartile range (IQR) of this data?

What is the standard deviation for this data?

Which is the better measure of dispersion? Why?

A. IQR. It is easier to find.

B. IQR. The boxplot shows outlier(s) and IQR is a resistant measure of dispersion.

C Standard deviation. It goes with mean which makes it better.

D. Standard deviation. The boxplot shows no outlier(s) so it is better to use in this case.
Answer #1

A. Mean is the better measure of central tendency because it consider the all data values B. Q1=1st quartile =median for the

D. Mean = x/n= 138/24 =5.75 Standard deviation s= /(x- ² v n-1 234.5 -=4.8207 24-1 E. Standard deviation is the better measur